OpenStudy (anonymous):

Two forces of 250 N and 575 N act on an object. The angle between them is 60 degrees. Determine the magnitude and the direction of the resultant force relative to the 250 N force.

OpenStudy (anonymous):



OpenStudy (anonymous):

So how would I start solving the problem?

OpenStudy (anonymous):

OK, I'm going to guess. Will it be as simple as adding two vectors?

OpenStudy (anonymous):

So,  F1 = 250 N F2 = 575 N f1x = 250N f1Y = 0 F2x = 575cos60 f2y = 575sin60 fx = (250N + 575cos60) = 537.5N [E] fy = (0+ 575sin60) = 497.96 N [N] angle = arc tan (497.96/537.5) fnet = Sqrt*((fx^2 +fy^2)) = 732.72 N [41.8 N of E]